OpenStudy (anonymous):

A rectangle has a width of 3 cm and a perimeter of 16 cm. What is the width of a similar rectangle with a length of 20 cm?

OpenStudy (anonymous):

First rect has length, 5, so new rect has 4* the length, thereby 4* the width. So, 4*3 = 12, is the width of new rectangle...

OpenStudy (anonymous):

first rect has a lenght of 13 cos perimeter=l+b since the rect are similar the then the ratios wud b similar  using ratios 20/13=x/3 x=4.6 p=l+b 20+4.6 24.6